a lamborghini can accelearate from rest to 28m/s in 2.4 s. assuming constant acceleration l, how fast will the lamborghini be going after 100 m?
Answer #1

where, acceleration, v2= Vitat 28 = o tax 2.4 a = 11.67 m/s2 speed after s=loom, 3 vit Vizo t=2.45 V2 = 28 m/svo? = 0+2x11.67 xlo |V3=48.3osmis